OpenStudy (secret-ninja):

Medal. Simplify the following fraction:

OpenStudy (secret-ninja):

\[\frac{ 16 }{ 14.5 }\]

OpenStudy (emmamink):

Let's find a common divisor between the two.

OpenStudy (secret-ninja):

I'm not sure how, since the denominator has a decimal, but I guess we can use the number 2?

ganeshie8 (ganeshie8):

do u have options ? ("simplify" is a very ambigious word)

OpenStudy (swissgirl):

Well you can always convert 14.5 into a fraction and then its simple

OpenStudy (emmamink):

Let's start by dividing by 2.

OpenStudy (secret-ninja):

I am trying to find slope, and when I used the slop formula I got this. I just need to make it smaller now. :p

OpenStudy (emmamink):

8/7.25.

OpenStudy (secret-ninja):

Dividing by 2 get 8/7.25

OpenStudy (secret-ninja):

Yep.

OpenStudy (emmamink):

1 3/4

OpenStudy (emmamink):

And there you have it!

OpenStudy (emmamink):

1 and 3/4, as I should do it

OpenStudy (emmamink):

11/4!

OpenStudy (secret-ninja):

So the slope is 1.75?

OpenStudy (secret-ninja):

Wait where did you get 11/4?

OpenStudy (emmamink):

I think....although I could be wrong.

OpenStudy (emmamink):

Yes, I should think so.

OpenStudy (emmamink):

1.75 as a fraction is 11/4

OpenStudy (secret-ninja):

Oh wait thats not 11 its 1 and 1/4

OpenStudy (skullpatrol):

$$\Huge \frac{16}{14.5}=\frac{16}{14 \frac{1}{2}}$$

OpenStudy (emmamink):

No I messed up there

OpenStudy (emmamink):

Wait.....the final answer is 7/4

OpenStudy (secret-ninja):

I'm so confused >.<

OpenStudy (emmamink):

Sorry about the confusion. Converted into decimal it would 1.75 or 7/4

OpenStudy (anonymous):

\[\frac{16}{14.5}=\frac{160}{145}=\frac{32}{29} \]

OpenStudy (emmamink):

There you have it!!!

OpenStudy (jhannybean):

There are several SEVERAL ways one can "simplify" the fraction, even using logarithmic rules, oh my goodness~

OpenStudy (secret-ninja):

Oh.. okay thanks. :]

OpenStudy (emmamink):

Do I get a medal?

OpenStudy (jhannybean):

@robtobey's method is probably the easiest, imo.

OpenStudy (secret-ninja):

I already gave you one >.<

OpenStudy (emmamink):

U did?

OpenStudy (secret-ninja):

Yep. Just reload ur page.

OpenStudy (emmamink):

Still dont have one

OpenStudy (emmamink):

Did you click best reponse?

OpenStudy (emmamink):

Ok!

OpenStudy (secret-ninja):

I gave it to you.

OpenStudy (emmamink):

I'm sure it will show up! Thanks a lot!

OpenStudy (secret-ninja):

Okie dokie bye

OpenStudy (secret-ninja):

No, thank you. ^-^

OpenStudy (emmamink):

Adioa!

ganeshie8 (ganeshie8):

who wants a medal

OpenStudy (emmamink):

I meant to say Adios

OpenStudy (emmamink):

I do!!!!

OpenStudy (emmamink):

Ya ur right. No medals people party over! Close it!

OpenStudy (emmamink):

And ur even a mod! U should b ashamed @ganeshie8

OpenStudy (secret-ninja):

It's been closed for several minutes now............

OpenStudy (emmamink):

ohhh..

OpenStudy (secret-ninja):

lol

OpenStudy (skullpatrol):

$$\Huge \frac{16}{14 \frac{1}{2}}= \frac{16}{ \frac{29}{2}}=\frac{32}{29} $$

OpenStudy (emmamink):

how did we get back to this?

OpenStudy (secret-ninja):

idk. He used latex, so it took a while. XD

OpenStudy (jhannybean):

@ganeshie8 my savior.

OpenStudy (emmamink):

Well @EmmaMink out guys! Peace!!!

ganeshie8 (ganeshie8):

No comments @EmmaMink

OpenStudy (emmamink):

I'm sorry!!! I have respect for mods. I did not mean to offend you

ganeshie8 (ganeshie8):

yeah i see my humor is bad today so i had to cut it short >.<

OpenStudy (secret-ninja):

lol. Well, not to be harsh to anyone, but the question is closed and I don't ant to have to keep coming back. Shows over, have a nice day ladies and gentlemen.